# Effective Fine-Tuning of Vision-Language Models for Accurate Galaxy Morphology Analysis

### v1

@@
Abstract
Galaxy morphology analysis involves classifying galaxies by their shapes and structures. For this task, directly training domain-specific models on large, annotated astronomical datasets is effective but costly. In contrast, fine-tuning vision foundation models on a smaller set of astronomical images is more resource-efficient but generally results in lower accuracy. To harness the benefits of both approaches and address their shortcomings, we propose GalaxAlign, a novel method that fine-tunes pre-trained foundation models to achieve high accuracy on astronomical tasks. Specifically, our method extends a contrastive learning architecture to align three types of data in fine-tuning: (1) a set of schematic symbols representing galaxy shapes and structures, (2) textual labels of these symbols, and (3) galaxy images. This way, GalaxAlign not only eliminates the need for expensive pretraining but also enhances the effectiveness of fine-tuning. Extensive experiments on galaxy classification and similarity search demonstrate that our method effectively fine-tunes general pre-trained models for astronomical tasks by incorporating domain-specific multi-modal knowledge.
1 Introduction
Galaxy morphology analysis involves studying galaxies based on their shapes and structures. This information is crucial for understanding galaxy formation and evolution, and it can be conveyed through natural language and schematic diagrams of galaxy images[29, 16]. As shown in Figure 1, schematic symbols paired with textual labels effectively capture the distinct characteristics of individual galaxies[3, 4]. These textual descriptions and schematic diagrams have proven useful in guiding amateur volunteers in galaxy image annotation[29]. In this paper, we explore how this multi-modal information can enhance foundation models for galaxy morphology analysis.
@@ 1 Introduction
Figure 1: Examples of galaxy images, corresponding schematic symbols and their textual labels. Images and textual labels are from the Galaxy10 DECaLS dataset [11], whereas schematic symbols are from the Galaxy Zoo 2 decision tree [7, 34].
Foundation models pre-trained on large-scale natural image datasets, such as ImageNet [21], perform well in various multimedia applications. However, researchers commonly believe that these pre-trained models are insufficient for astronomical images [31, 24, 27]. This concern arises because the domain-specific scientific data differ significantly from the general natural image datasets used for foundation models, leading to distribution shifts that reduce the effectiveness of these models when applied directly to galaxy analysis [27].
As a result, astronomical foundation models have typically relied on pretraining with large-scale astronomical datasets [16, 34, 7, 15]. These models are trained from scratch using extensive domain-specific datasets and subsequently fine-tuned for downstream tasks, overlooking the potential benefits of adapting publicly available vision foundation models that are pre-trained on natural datasets. Moreover, annotating galaxy images heavily depends on the efforts of experts and volunteers, requiring significant human resources and time [14, 29, 15]. Therefore, there is a pressing need for a method that reduces reliance on large domain-specific datasets and directly utilizes readily available pre-trained models trained on natural image data.
Based on the above observations, this work aims to investigate whether a galaxy model extended from pre-trained foundation models on natural image data can reduce the reliance on large astronomical datasets. To address this need, we introduce GalaxAlign, a tri-modal fine-tuning framework that adapts pre-trained CLIP models [20] for galaxy morphology analysis by integrating schematic symbols, textual descriptions, and galaxy images. Since human amateur volunteers can label astronomical images based on their knowledge learned from textual descriptions and schematic diagrams, we believe that these modalities can also assist models pre-trained on general datasets in performing classification tasks on astronomical images.
@@
  • Our method effectively extends models pre-trained on natural images for astronomical tasks, eliminating the need for large-scale astronomical datasets.
  • We reduce reliance on manually labeled data, providing a novel solution for galaxy morphology tasks.
  • We present a multi-modal learning architecture that incorporates text and symbols to adapt models pre-trained on general image datasets for use with galaxy images.
2 Background and Related Work
2.1 Foundation Models in Astrophysics
Unlike natural images, astrophysics images tend to have the following properties [27]:
@@ 2.1 Foundation Models in Astrophysics
Due to these differences, astronomers regard general vision models trained on natural images as inadequate for astronomical tasks. As such, instead of fine-tuning existing vision foundation models to fit astronomical data, Walmsley et al. proposed a galaxy morphology foundation model pre-trained on large-scale galaxy morphology datasets [31]. However, constructing such datasets is time-consuming and labor-intensive. For instance, the GZD-5 project, which classified 262,000 galaxies, spanned over three years from March 2017 to October 2020 [30]. Similarly, in the recent Galaxy Zoo campaign, 38,949 volunteers annotated a total of 105,000 galaxies over nearly two years (November 2020 to October 2022) [33].
To reduce the amount of time and labor required for creating large, labeled astronomical datasets, an alternative approach is to adapt existing foundation models pre-trained on natural images to astronomical tasks. Some representative pre-trained general vision foundation models including DINOv2 [19], MAE [10], MSN [1], supervised ResNet [9], and supervised ViT [8], have been fine-tuned to adapt to galaxy morphology tasks. However, their performance is generally poor [27]. These results suggest that adapting vision foundation models for astrophysics applications requires further considerations of data characteristics and alignment with domain-specific knowledge. Our work contributes to this effort.
Figure 2: Multi-modal annotation instructions for volunteers [35].
2.2 Citizen Science and Multi-Modal Data
Citizen science projects are the primary method for large-scale galaxy morphology annotation. In the Galaxy Zoo projects [16, 34, 30], volunteers annotate astronomical images based on specific instructions. This process heavily relies on providing amateur volunteers with simplified schematic symbols and natural language descriptions to guide them through the classification process.
Figure 2 illustrates an example of the annotation guidance available to volunteers on the Galaxy Zoo online platform [35]. Volunteers are presented with symbols that depict different morphological features, such as spiral arms, bars, or mergers, accompanied by concise textual descriptions of these features. In this annotation process, the combination of visual symbols and natural language descriptions helps non-experts connect their common sense with astronomical knowledge, enabling them to effectively complete labeling tasks.
@@ 2.2 Citizen Science and Multi-Modal Data
2.3 Multi-Modal Contrastive Learning Models
Vision-language models, such as CLIP [20], utilize contrastive learning to align visual and textual embeddings. These models obtain strong generalization capabilities by training on large-scale datasets that pair images with corresponding textual descriptions. Bowles et al. [3, 4] proposed a method to connect radio galaxy images with human-generated natural language descriptions to derive semantic morphology classes for classification, demonstrating that textual descriptions can align with distinct identifying features in the images [18].
Moreover, multi-modal contrastive learning has been widely applied in various scientific domains [17, 13, 22, 18, 15, 26], learning semantic representations effectively across diverse modalities. In this paper, we present the first application of associating astronomical data with three modalities—images, text, and schematic symbols—demonstrating that contrastive learning can effectively align these modalities.
3 Method
Figure 3: Overview of our GalaxAlign. Stage 1 is the initial finetuning with schematic symbols and images sharing a single encoder. Stage 2 is the finetuning with three separate encoders.
We introduce GalaxAlign, a tri-modal learning framework based on the CLIP architecture, for galaxy morphology tasks by incorporating and aligning three modalities: textual descriptions, astronomical images, and schematic symbols. Our approach contains three encoders: a text encoder for textual labels, an image encoder for galaxy images, and a symbol encoder for schematic symbols. Our finetuning is conducted in two stages: (1) Symbol-Image with Text Training and (2) Tri-modal Joint Training.
@@ 3 Method
  • Image Encoder: Initially it encodes both galaxy images and schematic symbols. In the second stage, the image encoder encodes galaxy images only.
@@
3.1 Textual Description
Our input text follows the common format A picture of a/an {class name}. Since the class name clearly and concisely describes the structure of the galaxies, it provides an efficient and identifying textual label. For example, sentences such as “A picture of an unbarred-spiral galaxy” or “A picture of a cigar round smooth galaxy” effectively capture the structural information of the galaxies.
3.2 Stage 1: Symbol-Image with Text Training
In the first stage, the Image Encoder takes both galaxy images and symbols as its input, whereas the Text Encoder processes corresponding text descriptions. This stage aligns the representations of images and symbols with text descriptions, setting up a shared multi-modal embedding space. The shared Image Encoder extracts patterns in galaxy shapes and structures that are consistent across symbols and images. Both encoders are initialized with pre-trained weights of the CLIP model [12] on natural images.
@@ 3.2 Stage 1: Symbol-Image with Text Training
3.3 Stage 2: Tri-Modal Joint Training
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Platform
@@
4.1.2 Datasets
In our experiments, we evaluate our method on two representative public galaxy datasets: (1) Galaxy10 DECaLS [11] and (2) GalaxyMNIST [28]. Galaxy10 contains 17,736 colored galaxy images divided in 10 classes, with each image of a size $256\times 256$ pixels. This dataset is from the DESI Legacy Imaging Surveys [6], which merges data from the Beijing-Arizona Sky Survey (BASS) [36], the DECam Legacy Survey (DECaLS) [2], and the Mayall z-band Legacy Survey [23]. In comparison, GalaxyMNIST [28], derived from Galaxy Zoo DECaLS [30], contains 10,000 galaxy images ($64\times 64$) of four morphological classes.
4.1.3 Evaluation Metrics
In this study, we evaluate the performance of GalaxAlign on classification tasks using Accuracy and F1 Score (macro). For the similarity search task, we use mean Average Precision (mAP), which reflects the ranking quality by calculating the average precision at various recall levels for each query and then averaging across all queries.
(a) Galaxy embedding visualization of Galaxy10 dataset.
(b) Galaxy embedding visualization of GalaxyMNIST dataset.
Figure 4: The t-SNE visualization for features extracted using Zoobot (MaxViT and ConvNeXT) and GalaxAlign (ViT-16 and ConvNeXT) on Galaxy10 and GalaxyMNIST datasets.
Figure 5: Visualisation of the representations learned by our method, illustrating similar galaxies occupying nearby regions in the feature space. This visualization is created using PCA to compress the representation to 2D and placing galaxy thumbnails at the locations of their corresponding galaxies in the grid.
4.1.4 Methods under Comparison
We evaluate GalaxAlign with the state-of-the-art foundation models, including mainstream general vision models and specialized astronomical foundation models trained on large-scale domain-specific datasets. The baseline models include:
  • MAE [10], DINOv2 [19], MSN [1]: Vision foundation models using self-supervised pretraining techniques. Following Lastufka et al. [27], we fine-tune the models using galaxy datasets of variant sizes in downstream tasks.
  • ViT-16 [8], ResNet-50 and ResNet-18 [9]: Standard vision models pretrained on ImageNet-1k [9] with a supervised method for classification.
  • Zoobot [32]: State-of-the-art astronomical models, pretrained specifically on large scale galaxy data from scratch. We selected MaxVIT and ConvNeXT as the backbone networks for Zoobot because they performed the best among all backbones [32].
Table 1: Models used in our comparison study.
@@ 4.1.4 Methods under Comparison
MAE [10] ViT-Base/16 ImageNet-1k [5]
DINOv2 [19] ViT-Base/14 LVD-142M [19]
MSN [1] ViT-Base/16 ImageNet-1k [5]
ViT-16 [8] ViT-Base/16 ImageNet-1k [5]
ResNet 50 [9] ResNet 50 ImageNet-1k [5]
ResNet 18 [9] ResNet 18 ImageNet-1k [5]
Zoobot [32] MaxViT-Base GZ DECaLS GZD-5 [30]
ConvNeXT-Base GZ DECaLS GZD-5 [30]
4.2 Feature Projections
To demonstrate GalaxAlign’s strong performance without relying on extensive domain-specific datasets, We present the embedding visualization comparing our method with Zoobot, which has been pretrained on large astronomical datasets and then fine-tuned on smaller datasets. Figure 4 provides a t-SNE visualization [25] of galaxy data embeddings learned by different models. Feature embeddings of other baseline methods are presented in the appendix.
In Figure 4 (a), both Zoobot models (MaxViT and ConvNeXT) display notable overlaps between complex classes including Barred Spiral Galaxy, Unbarred Tight Spiral and Unbarred Loose Spiral, indicating limited effectiveness in distinguishing fine morphological details. In contrast, our GalaxAlign models (ViT-16 and ConvNeXT) exhibit well-separated, clearly defined clusters with minimal overlap, demonstrating effectiveness in category distinction and feature representation. Meanwhile, GalaxAlign consistently maintains distinct clusters across all categories, reflecting its ability to capture essential morphological features accurately. In Figure 4 (b), Zoobot models exhibit some clustering but display considerable overlap in the Smooth Cigar and Edge-on Disk categories. In contrast, GalaxAlign models achieve more distinct, well-separated clusters across all categories. Overall, these results highlight GalaxAlign’s advantage over Zoobot in achieving compact intra-class clustering and distinct inter-class separability, underscoring its effectiveness in representing galaxy morphology.
Figure 5 shows a 2D grid visualization of galaxy images sampled from the GalaxyMNIST dataset, where high-dimensional features are reduced using PCA and aligned to a grid. The clusters and smooth transitions between similar galaxy types in the grid suggest that our model effectively encodes structural information in the feature space, allowing visually similar galaxies to occupy nearby regions. The visualization of images in the Galaxy10 dataset is presented in the appendix.
Table 2: Comparison of Classification Performance (mean and standard deviation) on GalaxyMNIST and Galaxy10

### v2

# GalaxAlign: Mimicking Citizen Scientists’ Multimodal Guidance for Galaxy Morphology Analysis

@@
Abstract.
Galaxy morphology analysis involves studying galaxies based on their shapes and structures. For such studies, fundamental tasks include identifying and classifying galaxies in astronomical images, as well as retrieving visually or structurally similar galaxies through similarity search. Existing methods either directly train domain-specific foundation models on large, annotated datasets or fine-tune vision foundation models on a smaller set of images. The former is effective but costly, while the latter is more resource-efficient but often yields lower accuracy. To address these challenges, we introduce GalaxAlign, a multimodal approach inspired by how citizen scientists identify galaxies in astronomical images by following textual descriptions and matching schematic symbols. Specifically, GalaxAlign employs a tri-modal alignment framework to align three types of data during fine-tuning: (1) schematic symbols representing galaxy shapes and structures, (2) textual labels for these symbols, and (3) galaxy images. By incorporating multimodal instructions, GalaxAlign eliminates the need for expensive pretraining and enhances the effectiveness of fine-tuning. Experiments on galaxy classification and similarity search demonstrate that our method effectively fine-tunes general pre-trained models for astronomical tasks by incorporating domain-specific multi-modal knowledge. Code is available at https://github.com/RapidsAtHKUST/GalaxAlign.
Multimodal Learning, Galaxy Morphology, Citizen Science
submissionid: 4688journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755600isbn: 979-8-4007-2035-2/2025/10ccs: Applied computing Astronomy
1. Introduction
Galaxy morphology analysis involves studying galaxies based on their shapes and structures. This information is crucial for understanding galaxy formation and evolution, and it can be conveyed through natural language and schematic diagrams of galaxy images(Walmsley et al., 2020; Lintott et al., 2011). As shown in Figure 1, schematic symbols paired with textual labels effectively capture the distinct characteristics of individual galaxies(Bowles et al., 2022, 2023). These textual descriptions and schematic diagrams have proven useful in instructing citizen scientists in galaxy image annotation(Walmsley et al., 2020), where the annotations from citizen scientists serve as a critical pipeline for processing massive astronomical datasets. In this paper, we explore how to train an AI-based citizen scientist for galaxy image annotation, demonstrating the use of multi-modal instructions to enhance existing foundation models for galaxy morphology analysis.
@@ 1. Introduction
Figure 1. Examples of galaxy images, corresponding schematic symbols and their textual labels. Images and textual labels are from the Galaxy10 DECaLS dataset (Henry, 2021), whereas schematic symbols are from the Galaxy Zoo 2 decision tree (Dieleman et al., 2015; Willett et al., 2013).
Foundation models pre-trained on large-scale natural image datasets, such as ImageNet (Russakovsky et al., 2015), perform well in various domain applications. However, researchers commonly believe that these pre-trained models are insufficient for astronomical images (Walmsley et al., 2022b; Slijepcevic et al., 2024; Voloshynovskyy, 2024). This concern arises because the domain-specific scientific data differ significantly from the general natural image datasets used for foundation models, leading to distribution shifts that reduce the effectiveness of these models when applied directly to galaxy analysis (Voloshynovskyy, 2024).
As a result, astronomical foundation models have typically relied on pretraining with large-scale astronomical datasets (Lintott et al., 2011; Willett et al., 2013; Dieleman et al., 2015; Lanusse et al., 2023). These models are trained from scratch using extensive domain-specific datasets and subsequently fine-tuned for downstream tasks, overlooking the potential benefits of adapting publicly available vision foundation models that are pre-trained on natural datasets. Moreover, annotating galaxy images heavily depends on the efforts of experts and volunteers, requiring significant human resources and time (Jiménez et al., 2023; Walmsley et al., 2020; Lanusse et al., 2023). Therefore, there is a pressing need for a method that reduces reliance on large domain-specific datasets and directly utilizes readily available pre-trained models trained on natural image data.
To address this need, we introduce GalaxAlign, a tri-modal framework that adapts pre-trained CLIP models (Radford et al., 2021) for galaxy morphology analysis by integrating schematic symbols, textual descriptions, and galaxy images. Since human amateur volunteers can label astronomical images based on their knowledge learned from textual descriptions and schematic diagrams, we believe that these modalities can also assist models pre-trained on general datasets in performing classification tasks on astronomical images.
@@
  • We propose GalaxAlign, the first tri-modal alignment framework that integrates galaxy images, schematic symbols, and natural language descriptions, effectively mimicking the cognitive strategy of citizen scientists in galaxy morphology analysis.
  • We design a two-stage multimodal adaptation pipeline that first unifies visual-symbolic representations through a shared encoder, and then transitions to modality-specific encoders for fine-grained alignment, enabling efficient and accurate transfer from general vision-language models to domain-specific astronomical tasks.
2. Background and Related Work
2.1. Foundation Models in Astrophysics
Unlike natural images, astrophysics images have the following properties (Voloshynovskyy, 2024):
@@ 2.1. Foundation Models in Astrophysics
Due to these differences, astronomers regard general vision models trained on natural images as inadequate for astronomical tasks. As such, instead of fine-tuning existing vision foundation models to fit astronomical data, Walmsley et al. proposed a galaxy morphology foundation model pre-trained on large-scale galaxy morphology datasets (Walmsley et al., 2022b). However, constructing such datasets is time-consuming and labor-intensive. For instance, the GZD-5 project, which classified 262,000 galaxies, spanned over three years from March 2017 to October 2020 (Walmsley et al., 2022a). Similarly, in the recent Galaxy Zoo campaign, 38,949 volunteers annotated a total of 105,000 galaxies over nearly two years (November 2020 to October 2022) (Walmsley et al., 2023b).
To reduce the amount of time and labor required for creating large, labeled astronomical datasets, an alternative approach is to adapt existing foundation models pre-trained on natural images to astronomical tasks. Some representative pre-trained general vision foundation models including DINOv2 (Oquab et al., 2023), MAE (He et al., 2022), MSN (Assran et al., 2022), supervised ResNet (He et al., 2016), and supervised ViT (Dosovitskiy, 2020), have been fine-tuned to adapt to galaxy morphology tasks. However, their performance is generally poor (Voloshynovskyy, 2024). These results suggest that adapting vision foundation models for astrophysics applications requires further considerations of data characteristics and alignment with domain-specific knowledge. Our work contributes to this effort.
Figure 2. Multi-modal annotation instructions for volunteers (Zoo, 2024).
2.2. Citizen Science and Multi-Modal Data
Citizen science projects are the primary method for large-scale galaxy morphology annotation. In the Galaxy Zoo projects (Lintott et al., 2011; Willett et al., 2013; Walmsley et al., 2022a), volunteers annotate astronomical images based on specific instructions. This process heavily relies on providing amateur volunteers with simplified schematic symbols and natural language descriptions to guide them through the classification process.
Figure 2 illustrates an example of the annotation guidance available to volunteers on the Galaxy Zoo online platform (Zoo, 2024). Volunteers are presented with symbols that depict different morphological features, such as spiral arms, bars, or mergers, accompanied by concise textual descriptions of these features. In this annotation process, the combination of visual symbols and natural language descriptions helps non-experts connect their common sense with astronomical knowledge, enabling them to effectively complete labeling tasks.
@@ 2.2. Citizen Science and Multi-Modal Data
2.3. Multi-Modal Contrastive Learning Models
Vision-language models, such as CLIP (Radford et al., 2021), utilize contrastive learning to align visual and textual embeddings. These models obtain strong generalization capabilities by training on large-scale datasets that pair images with corresponding textual descriptions. Bowles et al. (Bowles et al., 2022, 2023) proposed a method to connect radio galaxy images with human-generated natural language descriptions to derive semantic morphology classes for classification, demonstrating that textual descriptions aligns with distinct identifying features in the images.
Moreover, multi-modal contrastive learning has been widely applied in various scientific domains (Liu et al., 2023a; Imam et al., 2024; Sanchez-Fernandez et al., 2023; Mishra-Sharma et al., 2024; Lanusse et al., 2023; Vivanco Cepeda et al., 2024), learning semantic representations effectively across diverse modalities. In this paper, we present the first application of associating astronomical data with three modalities—images, text, and schematic symbols—demonstrating that contrastive learning can effectively align these modalities.
3. Method
Figure 3. Overview of our GalaxAlign. Stage 1 is the finetuning with schematic symbols and images sharing a single encoder. Stage 2 is the finetuning with 3 separate encoders.
We introduce GalaxAlign, a tri-modal learning framework based on the CLIP architecture, for galaxy morphology tasks by incorporating and aligning three modalities: textual descriptions, astronomical images, and schematic symbols.
@@ 3. Method
  • Image Encoder: Initially it encodes both galaxy images and schematic symbols to enable the learning of a shared representation of the two modalities. After the parameters from the shared representation are learned, the image encoder encodes galaxy images only.
@@
Our finetuning is conducted in two stages:
  • Stage 1: Symbol-image with text training.
  • Stage 2: Tri-modal joint training.
3.1. Textual Description
Our input text follows the common format A picture of a/an {class name}. Since the class name clearly and concisely describes the structure of the galaxies, it provides a succinct and distinguishing textual label. For example, sentences such as “A picture of an unbarred-spiral galaxy” or “A picture of a cigar round smooth galaxy” effectively capture the structural information of the galaxies.
3.2. Stage 1: Symbol-Image with Text Training
In the first stage, the Image Encoder takes both galaxy images and symbols as its input, whereas the Text Encoder processes corresponding text descriptions. This stage aligns the representations of images and symbols with text descriptions, setting up a shared multi-modal embedding space. The shared Image Encoder extracts patterns in galaxy shapes and structures that are consistent across symbols and images. Both encoders are initialized with pre-trained weights of the CLIP model (Ilharco et al., 2021) on natural images.
@@ 3.2. Stage 1: Symbol-Image with Text Training
3.3. Stage 2: Tri-Modal Joint Training
@@
This two-stage process enables the model to adapt pretrained terrestrial models to the specialized tasks of galaxy morphology classification, retrieval, and similarity search, utilizing both textual descriptions and schematic visual symbols as complementary modalities. Our experiments include comparisons with the state-of-the-art pretrained models under full-data, few-shot and zero-shot settings, demonstrating the effectiveness of our multimodal strategy in fine-tuning general models without the need for large-scale astronomical pretraining.
4. Experiments
In this section, we first present the visualization of galaxy feature embeddings from our model in comparison with a number of existing methods. Then, we conduct comparative studies on the accuracy performance of two tasks - galaxy morphology classification and similarity. Finally, we perform few-shot tests and ablation studies. More details regarding the implementation, datasets and additional experimental results are reported in the appendix.
4.1. Experimental Setup
4.1.1. Platform
@@
4.1.2. Datasets
In our experiments, we evaluate our method on two representative public galaxy datasets: (1) Galaxy10 DECaLS (Henry, 2021) and (2) GalaxyMNIST (Walmsley, 2022). Galaxy10 contains 17,736 colored galaxy images divided in 10 classes, with each image of a size $256\times 256$ pixels. This dataset is from the DESI Legacy Imaging Surveys (Dey et al., 2019), which merges data from the Beijing-Arizona Sky Survey (BASS) (Zou et al., 2017), the DECam Legacy Survey (DECaLS) (Blum et al., 2016), and the Mayall z-band Legacy Survey (Silva et al., 2016). In comparison, GalaxyMNIST (Walmsley, 2022), derived from Galaxy Zoo DECaLS (Walmsley et al., 2022a), contains 10,000 galaxy images ($64\times 64$) of four morphological classes.
4.1.3. Evaluation Metrics
In this study, we evaluate the performance of GalaxAlign on classification tasks using Accuracy and F1 Score (macro), providing a general sense of the model’s classification capability across the dataset as well as offering a balanced view of performance across all classes. For the similarity search task, we use mean Average Precision (mAP), which reflects the ranking quality by calculating the average precision at various recall levels for each query and then averaging across all queries.
Figure 4. The t-SNE visualization for features extracted using Zoobot (MaxViT and ConvNeXT) and GalaxAlign (ViT-16 and ConvNeXT) on Galaxy10 dataset.
Figure 5. Visualisation of the representations learned by our method, illustrating similar galaxies occupying nearby regions in the feature space. This visualization is created using PCA to compress the representation to 2D and placing galaxy thumbnails at the locations of their corresponding galaxies in the grid.
4.1.4. Methods under Comparison
We evaluate GalaxAlign with state-of-the-art foundation models, including mainstream general vision models, vision-language models and specialized astronomical foundation models trained on large-scale domain-specific datasets. The baseline models include:
MAE (He et al., 2022), DINOv2 (Oquab et al., 2023), MSN (Assran et al., 2022): Vision foundation models using self-supervised pretraining techniques. Following Lastufka et al. (Voloshynovskyy, 2024), we fine-tune the models using galaxy datasets of variant sizes in downstream tasks.
ViT-16 (Dosovitskiy, 2020), ResNet-50 and ResNet-18 (He et al., 2016): Standard vision models pretrained on ImageNet-1k (He et al., 2016) with a supervised method for classification.
CLIP (Radford et al., 2021): Standard vision-language models pretrained on Data-Comp-1B (Gadre et al., 2023) and Laion-2b (Schuhmann et al., 2022).
Zoobot (Walmsley et al., 2023a): State-of-the-art astronomical models, pretrained specifically on large-scale galaxy data from scratch. We selected MaxVIT and ConvNeXT as the backbone networks for Zoobot because they performed the best among all backbones (Walmsley et al., 2023a).
Table 1. Models used in our comparison study.
@@ 4.1.4. Methods under Comparison
MAE (He et al., 2022) ViT-Base/16 ImageNet-1k (Deng et al., 2009)
DINOv2 (Oquab et al., 2023) ViT-Base/14 LVD-142M (Oquab et al., 2023)
MSN (Assran et al., 2022) ViT-Base/16 ImageNet-1k (Deng et al., 2009)
ViT-16 (Dosovitskiy, 2020) ViT-Base/16 ImageNet-1k (Deng et al., 2009)
ResNet 50 (He et al., 2016) ResNet 50 ImageNet-1k (Deng et al., 2009)
ResNet 18 (He et al., 2016) ResNet 18 ImageNet-1k (Deng et al., 2009)
CLIP (Radford et al., 2021) ViT-16-Base DataComp-1B (Gadre et al., 2023)
ConvNext-Base LAION-2B (Schuhmann et al., 2022)
Zoobot (Walmsley et al., 2023a) MaxViT-Base GZ DECaLS GZD-5 (Walmsley et al., 2022a)
ConvNeXT-Base GZ DECaLS GZD-5 (Walmsley et al., 2022a)
4.2. Feature Projections
To demonstrate GalaxAlign’s strong performance without relying on extensive domain-specific datasets, we present the embedding visualization comparing our method with Zoobot, which has been pretrained on large astronomical datasets and then fine-tuned on smaller datasets. Figure 4 provides a t-SNE visualization (Van der Maaten and Hinton, 2008) of galaxy data embeddings of Galaxy10 dataset learned by different models. Feature embeddings of other baseline methods and datasets are presented in the appendix.
In Figure 4 (a), both Zoobot models (MaxViT and ConvNeXT) display notable overlaps between complex classes including Barred Spiral Galaxy, Unbarred Tight Spiral and Unbarred Loose Spiral, indicating limited performance in distinguishing fine morphological details. In contrast, our GalaxAlign models (ViT-16 and ConvNeXT) exhibit well-separated, clearly defined clusters with minimal overlap, demonstrating effectiveness in category distinction and feature representation. Meanwhile, GalaxAlign consistently maintains distinct clusters across all categories, reflecting its ability to capture essential morphological features accurately. In Figure 4 (b), Zoobot models exhibit some clustering but display considerable overlap in the Smooth Cigar and Edge-on Disk categories. In contrast, GalaxAlign models achieve more distinct, well-separated clusters across all categories. Overall, these results highlight GalaxAlign’s advantage over Zoobot in achieving compact intra-class clustering and distinct inter-class separability, underscoring its effectiveness in representing galaxy morphology.
Figure 5 shows a 2D grid visualization of galaxy images sampled from the GalaxyMNIST dataset, where high-dimensional features are reduced using PCA and aligned to a grid. The clusters and smooth transitions between similar galaxy types in the grid suggest that our model effectively encodes structural information in the feature space, allowing visually similar galaxies to be placed in close proximity. The visualization of images in the Galaxy10 dataset is presented in the appendix.
Table 2. Comparison of Classification Performance (mean and standard deviation) on GalaxyMNIST and Galaxy10